# Survey on Publicly Available Sinhala Natural Language Processing Tools and Research

### v21

@@ 3.10 Text Classification
Text classification is a popular application on the semantic layer of the NLP stack. A very basic Sinhala text classification using Naïve Bayes Classifier, Zipf’s Law Behavior, and SVMs was attempted by Gallege [198]. A smaller implementation of Sinhala news classification has been attempted by de Silva [23]. As mentioned in Section 3.2, their news corpus is publicly available [^82] https://osf.io/tdb84/. Another attempt at Sinhala text classification using six popular rule-based algorithms was done by Lakmali and Haddela [199]. Even though they talk about building a corpus named SinNG5, they do not indicate of means for others to obtain the said corpus. Another study by Kumari and Haddela [200] utilizes the SinNG5 corpus as the data set for their attempt to use LIME [201] for human interpretability of Sinhala document classification. However, they too do not provide access corpus. Nanayakkara and Ranathunga [202] have implemented a system which uses corpus-based similarity measures for Sinhala text classification.  Gunasekara and Haddela [203, 204] claim to have created a context-aware stop word extraction method for Sinhala text classification based on simple TF-IDF. An LSTM based textual entailment system for Sinhala was proposed by Jayasinghe and Sirts [205]Demotte and Ranathunga [206] proposed a dual-state capsule network architecture for text classification where they demonstrated their methodology on the Sinhala data set established by Senevirathne et al. [207]. The attempt by Sameemdeen and Selvanthan [208] considers three classical machine learning algorithms (Naïve Bayes,SVM,and KNN) and then goes on to briefly discusses the pros and cons of previous attempts by: Nanayakkara and Ranathunga [202]Gunasekara and Haddela [203]Lakmali and Haddela [199], and Buddhika et al. [209]. Then they propose active learning [210, 211] as an alternative. However, no experimental results of how active learning would improve Sinhala text classification are given. The study by Bandara et al. [212] proposed and ontology-based approach for Sinhala fake news detection. However, their literature survey did not cover seminal papers in OBIE such as the work by Wimalasuriya and Dou [26]. This has impacted the overall methodology that was presented. A novel, domain-idependant, and domain-adaptive text classification framework named AdaptText for Sinhala text classification was proposed by Kodithuwakku and Hettiarachchi [213]. A simple TF-IDF based Sinhala text classification system was proposed by Koralage [214]. The study by [215] used Genetic Algorithm on Lucene [^83] https://lucene.apache.org/ search queries to obtain interpretable classification models for Sinhala text documents. The approach proposed by Rathnayake et al. [216] uses adapter-based [217, 218, 219, 220, 221, 222] fine-tuning on XLM-R [223], for code-mixed and code-switched text. Kirindage and Godewithana [224] used LDA [225] to develop Sinhala news topic hierarchies and categorize Sinhala news documents using the said topic hierarchies. Hettigoda [226] collected English-Sinhala code-mixed comments from Facebook pages belonging to Clothing industrial online businesses in Sri Lanka and classified them in to the classes: Inquiries, Maybe Inquiries, and Not Inquiries. A comparative analysis on BERT based models for Sinhala text classification was conducted by Dhananjaya et al. [98]. Both their code [^84] https://github.com/nlpcuom/Sinhala-text-classification and data (sin-cc-15M corpus8585853https://tinyurl.com/42un7a9y, Sinhala News source classification data set [^86] https://bit.ly/3PBTyfo, Sinhala News category classification data set [^87] https://bit.ly/3AsfK7u, and Sinhala writing style classification data set [^88] https://bit.ly/3PARuUS) are publicly available. The study by Wijayarathna and Jayalal [227] used classical machine learning techniques including Random Forest to classify fake news in Sinhala on Twitter. Chathuranga and Ranathunga [228] used capsule-based methods recommended by Senevirathne et al. [207] to classify Sinhala-English code-mixed data. The work by Weerasiri et al. [229] compared Word2Vec, FastText and Doc2vec [230] against TF-IDF for Sinhala news document classification. The study by Caldera et al. [231] used stacked LSTM to classify Sinhala and Singlish text discussing COVID-19 on Youtube and Twitter.
@@ 3.11 Sentiment Analysis
The work by Jayawickrama et al. [243, 244] used the data set released by Wijeratne and de Silva [90] to predict the reactions induced by Sinhala Facebook posts. They then extended the work [245] and compared the results obtained with their data set against that of Senevirathne et al. [207]Aththanayaka and Naleer [246] used Random forest, Support vector machines, and Multinomial Naïve Bayes models to analyse sentiment in Sinhala-English code-mixed text from social media. The ACTSEA dataset [^90] https://bit.ly/3oZoziI for Sinhala sentiment analysis was introduced by Jenarthanan et al. [124]. The dataset contains 318,308 Sinhala tweets annotated with emotions.
@@ 3.12 Hate Speech Detection
A machine learning approach to detect hate speech in Sinhala was proposed by De Silva [248]. A feature model and a data set [^91] https://bit.ly/3FTNaMZ for Sinhala hate speech detection for youtube was proposed by De Saa and Ranathunga [117]Sandaruwan et al. [249] have attempted to identify abusive Sinhala comments in social media using text mining and machine learning techniques. A cyberbullying comment classification study for Sinhala was conducted by [250] where they used classical machine learning algorithms. The study by Hettiarachchi et al. [251] used classical machine learning methods to detect hate speech in Romanized Sinhala social media posts. While the basic idea is the same, they have avoided mentioning transliteration in their paper. The study by Samarasinghe et al. [252] proposed using CNNs for detecting hate speech in Sinhala text. Kariyawasam [253] proposed a machine learning approach for identifying toxic Sinhala language on social media. Guruge et al. [254] used an ensemble of Naïve Bayes, Support Vector Machine, XGBoost, MLP, and AdaBoost to detect hate speech in 49019 Tweets they collected from February of 2021 to April of 2021. Sandaruwan et al. [255] collected 3000 Sinhala comments corpus and conducted Multinomial Naïve Bayes to detect hate speech. Sheran [256] claims to have used machine learning to detect hate speech written in Sinhala or Singlish on social media. A deep learning approach for detecting hate speech in Sinhala tweets was explored by Munasinghe and Thayasivam [257]. The study by Shalinda and Munasinghe [258] utilized classical machine learning techniques such as Linear Support vector classifier, Random Forest Classification, SGD classifier, Logistic Regression, XGBoost classifier and multinomial Naive Bayes classifier on both Sinhala and Singlish (Romanized Sinhala) to identify hate words. They report that the SGD classifier over TF-IDF with uni-grams and bi-grams gives them the highest accuracy. Gamage et al. [259] conducted a comparative analysis on a number of embedding systems as well as classical frequency-based methods for Sinhala hate speech detection. The study by Fernando et al. [260] also claims to use machine learning and deep learning to detect hate speech in Sinhala. The study by Perera et al. [261] predicted Sinhala hate speech using user behaviour on twitter by applying ensemble Machine Learning to classify them by the five big personality traits. The follow-up study by Perera et al. [118] analyses Sinhala hate speech propagation on Twitter. They have also published a dataset [^92] https://github.com/Isurie/Text-Classification-Module/tree/master/Dataset of 1600+ annotated Sinhala tweets. Rajapaksha et al. [262] used deep learning to identify trending periods of hate topics on Twitter in Sinhala. Arachchi et al. [263] used a web-based tool to translate between Sinhala and English to detect inappropriate word usage in Sinhala.
@@
After the release of the SOLD dataset by Ranasinghe et al. [128] for the hate speech detection task in Sinhala, a number of subsequent works have been conducted using it. Fernando and Deng [264] used traditional machine learning algorithms to predict Sinhala hate speech. Ojo et al. [265] used mBERT to identify hate and offensive content in Sinhala under HASOC 2023, Task 1A [^93] https://hasocfire.github.io/hasoc/2023/task1.html which included the SOLD dataset. Ranasinghe and Zampieri [266] compared the results of pre-trained mT5 on Sinhala tweets from the SOLD dataset against results for German, Spanish, Hindi, Korean, and Marathi. The work by Bestgen [267] compared the result of classical machine learning algorithms on simple n-gram features for the hate speech detection task on the SOLD dataset for Sinhala against results for Assamese and Bengali. The study by Narayan et al. [268] compared the effectiveness of a number of pre-trained deep learning models for Sinhala hate speech detection. They showed that XLM Roberta Base achieves a macro-f1 of 83.49% over the LSTM baseline of 75.30%.
@@ 3.13 Fake News Detection
A dataset consisting of Sinhala documents drawn from Sri Lankan news websites was published by Jayawickrama et al. [115] along with the benchmark misinformation classification models. A hybrid approach to detect Sinhala fake news on Social media was proposed by Wijayarathna and Jayalal [269] where the text content of the post is checked against credible sources and the authenticity of the user account posting the relevant post is evaluated by a rule-based points allocation schema. Wijayarathna and Jayalal [270] collected a set of 120 fake news tweets and 250 non-fake news tweets which they then converted to vectors by taking the fasttext vector for words and averaging. The vector representations of the tweets were then compared to predict whether the news containing it is real or fake. The study by Udurawana et al. [271] proposes to use an accuracy score (obtained by analysing the news content) and a credibility score (obtained by using a scoring mechanism) to detect fake news in Sinhala text. They also incorporate a module that classifies on the basis of passive aggressiveness. The study by Adihetti and Jayalal [272] used autoencoders to detect Sinhala fake news from social media posts.
@@ 3.14 Word Sense Disambiguation
There have been multiple attempts to do word sense disambiguation (WSD) [273, 274, 275, 276, 277] for Sinhala. For this, Arukgoda et al. [159] have proposed a system named Aruth based on the Lesk Algorithm[278]. An online tool [^94] http://aruth.herokuapp.com/, an API [^95] https://bit.ly/3sJEYbS of the algorithm, and code along with data on github [^96] https://github.com/jseanm1/aruthSWSD are available. For the same task, Marasinghe et al. [279] have proposed a system based on probabilistic modeling. A dialogue act recognition system which utilizes simple classification algorithms has been proposed by Palihakkara et al. [280]. A word sense disambiguation tool named Sinsense was introduced by Subasingha [281]. They used cross-lingual sense disambiguation where English sense disambiguation was used to obtain Sinhala sense disambiguation. However, neither their tool nor their full research is publicly available.
@@ 3.15 Text Summarizing
A deterministic process flow for automatic Sinhala text summarizing was proposed by Welgama [282]. The study by Wimalasuriya [283], which has the same name as the above work by Welgama [282], uses the graph-based TextRank algorithm for automatic Sinhala text summarizing. The use case of Sinhala Text summarization for government gazettes was explored by Jayawardane [284]. The study by Rathnayake et al. [285] compared the results of extractive and abstractive summarization on Sinhala textbooks. The study by Jahan and Wijesekara [286] compared the abstractive summarizing methods of TF-IDF and Text-Rank for Sinhala using ROUGE as the evaluation score. They concluded TF-IDF to be the superior choice. Hasan et al. [137] introduced the XL-sum [^97] https://github.com/csebuetnlp/xl-sum/tree/master dataset which includes 3,414 Sinhala documents and their summaries collected from Sinhala BBC website [^98] https://www.bbc.com/sinhala. The M3LS [^99] https://github.com/anubhav-jangra/M3LS multi-modal dataset introduced by Verma et al. [138] contains 10,148 Sinhala documents, relevant images, and their summaries collected from Sinhala BBC website100100100https://www.bbc.com/sinhala. Further, they claim that MT5 [287] obtains the best ROUGE-1 and ROUGE-L f scores for their Sinhala data set.
@@ 3.17 Phonological Tools
On the case of phonological layer, a report on Sinhala phonetics and phonology was published by Wasala and Gamage [289]Wickramasinghe et al. [290] discussed the practical issues in developing Sinhala Text-to-Speech and Speech Recognition systems. The Massively Multilingual Speech (MMS)101101101https://github.com/facebookresearch/fairseq/tree/main/examples/mms data set created by Pratap et al. [131] has Sinhala data for the Spoken Language Identification (LID) task.
@@ 3.17.1 Text-to-Speech
Based on the earlier work by Weerasinghe et al. [291]Wasala et al. [292] have developed methods for Sinhala grapheme-to-phoneme conversion along with a set of rules for schwa epenthesis. This work was then extended by Nadungodage et al. [293]Weerasinghe et al. [294] developed a Sinhala text-to-speech system. However, it is not publicly accessible. They internally extended it to create a system capable of helping a mute person achieve synthesized real-time interactive voice communication in Sinhala [295]. A rule based approach for automatic segmentation of a small set of Sinhala text into syllables was proposed by Kumara et al. [296]. An ew prosodic phrasing method to help with Sinhala Text-to-Speech process was proposed by Bandara et al. [297, 298, 299]Sodimana et al. [300] proposed a text normalization methodology for Sinhala text-to-speech systems. Further, Sodimana et al. [301] formalized a step-by-step process for building text-to-speech voices for Sinhala. Both Jayamanna [302] and Mishangi [303] have created Sinhala document readers for visually impaired persons to be used on Android devices. An OCR and Text-to-Speech system for Sinhala named Bhashitha was proposed by De Zoysa et al. [304]. The works by Lakmal et al. [305] and Senarathna et al. [306] adapted MaryTTS [307] to synthesize Sinhala speech. The study by Jayawardhana et al. [308] used Deep Voice [309] for Sinhala and English TTS. Gamage et al. [310] included a Sinhala text-to-speech module as one of the three modules present in their currency recognition system.  Anuradha and Thelijjagoda [311] proposed a machine translation system to convert Sinhala and English Braille documents into voice. A separate group has done work on Sinhala text-to-speech systems independent to above [312].

### v22

@@ 3.10 Text Classification
Text classification is a popular application on the semantic layer of the NLP stack. A very basic Sinhala text classification using Naïve Bayes Classifier, Zipf’s Law Behavior, and SVMs was attempted by Gallege [205]. A smaller implementation of Sinhala news classification has been attempted by de Silva [24]. As mentioned in Section 3.2, their news corpus is publicly available [^90] https://osf.io/tdb84/. Another attempt at Sinhala text classification using six popular rule-based algorithms was done by Lakmali and Haddela [206]. Even though they talk about building a corpus named SinNG5, they do not indicate of means for others to obtain the said corpus. Another study by Kumari and Haddela [207] utilizes the SinNG5 corpus as the data set for their attempt to use LIME [208] for human interpretability of Sinhala document classification. However, they too do not provide access corpus. Nanayakkara and Ranathunga [209] have implemented a system which uses corpus-based similarity measures for Sinhala text classification.  Gunasekara and Haddela [210, 211] claim to have created a context-aware stop word extraction method for Sinhala text classification based on simple TF-IDF. An LSTM based textual entailment system for Sinhala was proposed by Jayasinghe and Sirts [212]Demotte and Ranathunga [213] proposed a dual-state capsule network architecture for text classification where they demonstrated their methodology on the Sinhala data set established by Senevirathne et al. [214]. The attempt by Sameemdeen and Selvanthan [215] considers three classical machine learning algorithms (Naïve Bayes,SVM,and KNN) and then goes on to briefly discusses the pros and cons of previous attempts by: Nanayakkara and Ranathunga [209]Gunasekara and Haddela [210]Lakmali and Haddela [206], and Buddhika et al. [216]. Then they propose active learning [217, 218] as an alternative. However, no experimental results of how active learning would improve Sinhala text classification are given. The study by Bandara et al. [219] proposed and ontology-based approach for Sinhala fake news detection. However, their literature survey did not cover seminal papers in OBIE such as the work by Wimalasuriya and Dou [29]. This has impacted the overall methodology that was presented. A novel, domain-idependant, and domain-adaptive text classification framework named AdaptText for Sinhala text classification was proposed by Kodithuwakku and Hettiarachchi [220]. A simple TF-IDF based Sinhala text classification system was proposed by Koralage [221]. The study by [222] used Genetic Algorithm on Lucene [^91] https://lucene.apache.org/ search queries to obtain interpretable classification models for Sinhala text documents. The approach proposed by Rathnayake et al. [223] uses adapter-based [224, 225, 226, 227, 228, 229] fine-tuning on XLM-R [230], for code-mixed and code-switched text. Kirindage and Godewithana [231] used LDA [232] to develop Sinhala news topic hierarchies and categorize Sinhala news documents using the said topic hierarchies. Hettigoda [233] collected English-Sinhala code-mixed comments from Facebook pages belonging to Clothing industrial online businesses in Sri Lanka and classified them in to the classes: Inquiries, Maybe Inquiries, and Not Inquiries. A comparative analysis on BERT based models for Sinhala text classification was conducted by Dhananjaya et al. [101]. Both their code [^92] https://github.com/nlpcuom/Sinhala-text-classification and data (sin-cc-15M corpus9393933https://tinyurl.com/42un7a9y, Sinhala News source classification data set [^94] https://bit.ly/3PBTyfo, Sinhala News category classification data set [^95] https://bit.ly/3AsfK7u, and Sinhala writing style classification data set [^96] https://bit.ly/3PARuUS) are publicly available. The study by Wijayarathna and Jayalal [234] used classical machine learning techniques including Random Forest to classify fake news in Sinhala on Twitter. Chathuranga and Ranathunga [235] used capsule-based methods recommended by Senevirathne et al. [214] to classify Sinhala-English code-mixed data. The work by Weerasiri et al. [236] compared Word2Vec, FastText and Doc2vec [237] against TF-IDF for Sinhala news document classification. The study by Caldera et al. [238] used stacked LSTM to classify Sinhala and Singlish text discussing COVID-19 on Youtube and Twitter.
@@ 3.11 Sentiment Analysis
The work by Jayawickrama et al. [250, 251] used the data set released by Wijeratne and de Silva [93] to predict the reactions induced by Sinhala Facebook posts. They then extended the work [252] and compared the results obtained with their data set against that of Senevirathne et al. [214]Aththanayaka and Naleer [253] used Random forest, Support vector machines, and Multinomial Naïve Bayes models to analyse sentiment in Sinhala-English code-mixed text from social media. The ACTSEA dataset [^98] https://bit.ly/3oZoziI for Sinhala sentiment analysis was introduced by Jenarthanan et al. [128]. The dataset contains 318,308 Sinhala tweets annotated with emotions. The study by Dhananjaya et al. [254] uses a sentiment lexicon of a high‐resource language to fine-tune Pre‐trained multilingual language models (PMLMs) such as mBERT and XLM‐R on an intermediate task which in turn is then used on the sentiment classification task of Sinhala.
@@ 3.12 Hate Speech Detection
A machine learning approach to detect hate speech in Sinhala was proposed by De Silva [256]. A feature model and a data set [^99] https://bit.ly/3FTNaMZ for Sinhala hate speech detection for youtube was proposed by De Saa and Ranathunga [121]Sandaruwan et al. [257] have attempted to identify abusive Sinhala comments in social media using text mining and machine learning techniques. A cyberbullying comment classification study for Sinhala was conducted by [258] where they used classical machine learning algorithms. The study by Hettiarachchi et al. [259] used classical machine learning methods to detect hate speech in Romanized Sinhala social media posts. While the basic idea is the same, they have avoided mentioning transliteration in their paper. The study by Samarasinghe et al. [260] proposed using CNNs for detecting hate speech in Sinhala text. Kariyawasam [261] proposed a machine learning approach for identifying toxic Sinhala language on social media. Guruge et al. [262] used an ensemble of Naïve Bayes, Support Vector Machine, XGBoost, MLP, and AdaBoost to detect hate speech in 49019 Tweets they collected from February of 2021 to April of 2021. Sandaruwan et al. [263] collected 3000 Sinhala comments corpus and conducted Multinomial Naïve Bayes to detect hate speech. Sheran [264] claims to have used machine learning to detect hate speech written in Sinhala or Singlish on social media. A deep learning approach for detecting hate speech in Sinhala tweets was explored by Munasinghe and Thayasivam [265]. The study by Shalinda and Munasinghe [266] utilized classical machine learning techniques such as Linear Support vector classifier, Random Forest Classification, SGD classifier, Logistic Regression, XGBoost classifier and multinomial Naive Bayes classifier on both Sinhala and Singlish (Romanized Sinhala) to identify hate words. They report that the SGD classifier over TF-IDF with uni-grams and bi-grams gives them the highest accuracy. Gamage et al. [267] conducted a comparative analysis on a number of embedding systems as well as classical frequency-based methods for Sinhala hate speech detection. The study by Fernando et al. [268] also claims to use machine learning and deep learning to detect hate speech in Sinhala. The study by Perera et al. [269] predicted Sinhala hate speech using user behaviour on twitter by applying ensemble Machine Learning to classify them by the five big personality traits. The follow-up study by Perera et al. [122] analyses Sinhala hate speech propagation on Twitter. They have also published a dataset100100100https://github.com/Isurie/Text-Classification-Module/tree/master/Dataset of 1600+ annotated Sinhala tweets. Rajapaksha et al. [270] used deep learning to identify trending periods of hate topics on Twitter in Sinhala. Arachchi et al. [271] used a web-based tool to translate between Sinhala and English to detect inappropriate word usage in Sinhala. The study by Ehelepola [272] analyzed Sinhala text from social media and e-commerce sites for hate speech. The work by Dikwatta et al. [273] used supervised algorithms to detect Sinhala hate speech in text from image posts on the Internet (memes). Wickramaarachchi et al. [274] used LSTM on BART to compare the title and description of Sinhala youtube videos to their audio in order to determine if they contain hate speech.
@@
After the release of the SOLD dataset by Ranasinghe et al. [132] for the hate speech detection task in Sinhala, a number of subsequent works have been conducted using it. Fernando and Deng [275] used traditional machine learning algorithms to predict Sinhala hate speech. Ranasinghe and Zampieri [276] compared the results of pre-trained mT5 on Sinhala tweets from the SOLD dataset against results for German, Spanish, Hindi, Korean, and Marathi. The work by Bestgen [277] compared the result of classical machine learning algorithms on simple n-gram features for the hate speech detection task on the SOLD dataset for Sinhala against results for Assamese and Bengali. Ranasinghe et al. [278] used the SOLD dataset for the Hate Speech And Offensive Content (HASOC) identification task101101101https://hasocfire.github.io/hasoc/2023/task1.html they introduced. In completing the task, the study by Narayan et al. [279] compared the effectiveness of a number of pre-trained deep learning models for Sinhala hate speech detection. They showed that XLM Roberta Base achieves a macro-f1 of 83.49% over the LSTM baseline of 75.30%. For the same task, Ojo et al. [280] used mBERT to identify hate and offensive content in Sinhala.
@@ 3.13 Fake News Detection
A dataset consisting of Sinhala documents drawn from Sri Lankan news websites was published by Jayawickrama et al. [119] along with the benchmark misinformation classification models. A hybrid approach to detect Sinhala fake news on Social media was proposed by Wijayarathna and Jayalal [281] where the text content of the post is checked against credible sources and the authenticity of the user account posting the relevant post is evaluated by a rule-based points allocation schema. Wijayarathna and Jayalal [282] collected a set of 120 fake news tweets and 250 non-fake news tweets which they then converted to vectors by taking the fasttext vector for words and averaging. The vector representations of the tweets were then compared to predict whether the news containing it is real or fake. The study by Udurawana et al. [283] proposes to use an accuracy score (obtained by analysing the news content) and a credibility score (obtained by using a scoring mechanism) to detect fake news in Sinhala text. They also incorporate a module that classifies on the basis of passive aggressiveness. The study by Adihetti and Jayalal [284] used autoencoders to detect Sinhala fake news from social media posts. The study by Wickramaarachchi et al. [274] detected fake content in Sinhala youtube videos by comparing their title and description to the audio using LSTM on BART.
@@ 3.14 Word Sense Disambiguation
There have been multiple attempts to do word sense disambiguation (WSD) [285, 286, 287, 288, 289] for Sinhala. For this, Arukgoda et al. [166] have proposed a system named Aruth based on the Lesk Algorithm[290]. An online tool102102102http://aruth.herokuapp.com/, an API103103103https://bit.ly/3sJEYbS of the algorithm, and code along with data on github104104104https://github.com/jseanm1/aruthSWSD are available. For the same task, Marasinghe et al. [291] have proposed a system based on probabilistic modeling. A dialogue act recognition system which utilizes simple classification algorithms has been proposed by Palihakkara et al. [292]. A word sense disambiguation tool named Sinsense was introduced by Subasingha [293]. They used cross-lingual sense disambiguation where English sense disambiguation was used to obtain Sinhala sense disambiguation. However, neither their tool nor their full research is publicly available.
@@ 3.15 Text Summarizing
A deterministic process flow for automatic Sinhala text summarizing was proposed by Welgama [294]. The study by Wimalasuriya [295], which has the same name as the above work by Welgama [294], uses the graph-based TextRank algorithm for automatic Sinhala text summarizing. The use case of Sinhala Text summarization for government gazettes was explored by Jayawardane [296]. The study by Rathnayake et al. [297] compared the results of extractive and abstractive summarization on Sinhala textbooks. The study by Jahan and Wijesekara [298] compared the abstractive summarizing methods of TF-IDF and Text-Rank for Sinhala using ROUGE as the evaluation score. They concluded TF-IDF to be the superior choice. Hasan et al. [141] introduced the XL-sum105105105https://github.com/csebuetnlp/xl-sum/tree/master dataset which includes 3,414 Sinhala documents and their summaries collected from Sinhala BBC website106106106https://www.bbc.com/sinhala. The M3LS107107107https://github.com/anubhav-jangra/M3LS multi-modal dataset introduced by Verma et al. [142] contains 10,148 Sinhala documents, relevant images, and their summaries collected from Sinhala BBC website108108108https://www.bbc.com/sinhala. Further, they claim that MT5 [299] obtains the best ROUGE-1 and ROUGE-L f scores for their Sinhala data set.
@@ 3.17 Phonological Tools
On the case of phonological layer, a report on Sinhala phonetics and phonology was published by Wasala and Gamage [301]Wickramasinghe et al. [302] discussed the practical issues in developing Sinhala Text-to-Speech and Speech Recognition systems. The Massively Multilingual Speech (MMS)109109109https://github.com/facebookresearch/fairseq/tree/main/examples/mms data set created by Pratap et al. [135] has Sinhala data for the Spoken Language Identification (LID) task. The meta-study by Al-Fraihat et al. [303] compared the status of Sinhala speech recognition research against 17 other languages.
@@ 3.17.1 Text-to-Speech
Based on the earlier work by Weerasinghe et al. [304]Wasala et al. [305] have developed methods for Sinhala grapheme-to-phoneme conversion along with a set of rules for schwa epenthesis. This work was then extended by Nadungodage et al. [306]Weerasinghe et al. [307] developed a Sinhala text-to-speech system. However, it is not publicly accessible. They internally extended it to create a system capable of helping a mute person achieve synthesized real-time interactive voice communication in Sinhala [308]. A rule based approach for automatic segmentation of a small set of Sinhala text into syllables was proposed by Kumara et al. [309]. An ew prosodic phrasing method to help with Sinhala Text-to-Speech process was proposed by Bandara et al. [310, 311, 312]Sodimana et al. [313] proposed a text normalization methodology for Sinhala text-to-speech systems. Further, Sodimana et al. [314] formalized a step-by-step process for building text-to-speech voices for Sinhala. Both Jayamanna [315] and Mishangi [316] have created Sinhala document readers for visually impaired persons to be used on Android devices. An OCR and Text-to-Speech system for Sinhala named Bhashitha was proposed by De Zoysa et al. [317]. The works by Lakmal et al. [318] and Senarathna et al. [319] adapted MaryTTS [320] to synthesize Sinhala speech. The study by Jayawardhana et al. [321] used Deep Voice [322] for Sinhala and English TTS. Gamage et al. [323] included a Sinhala text-to-speech module as one of the three modules present in their currency recognition system. The study by Madhusha et al. [324] claims to have created a mobile app with Sinhala Text-to-Speech and OCR to read books for visually impaired students.  Anuradha and Thelijjagoda [325] proposed a machine translation system to convert Sinhala and English Braille documents into voice. A separate group has done work on Sinhala text-to-speech systems independent to above [326].